OpenStudy (lυἶცἶ0210):

Just wanted to make sure I was on the right track before continuing.. looks a bit messy

OpenStudy (lυἶცἶ0210):

\(\Large y'+y=tsint\) and y(0)=0 \(\Large [SF(s)-f(0)]+[F(s)]=\frac{2S}{(S^2+1)^2} \) \(\Large SF(s)+F(s)-0=\frac{2S}{(S^2+1)^2} \) \(\Large F(s)(S+1) = \frac{2S}{(S^2+1)^2}\) \( \Large F(s)=\frac{2S}{(S^2+1)^2(S+1)} \)

OpenStudy (anonymous):

It looks right